Qiang places the amount of $98 in a special bank account that guarantees to provide $188 in 5 years. What rate of return would Qiang earn on this opportunity?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

188=98*(1+r/100)^5

(188/98)^(1/5)=(1+r/100)

(1+r/100)=1.1392

r=1.1392-1

=13.92%(Approx).